OpenStudy (nonoww):

What is the equation in point−slope form of the line passing through (0, 6) and (1, 3)?

OpenStudy (nonoww):

Pleas help

TheSmartOne (thesmartone):

First find the slope \(\sf\Large Slope = \frac{\color{blue}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{purple}{x_1}}\) Where the two points are in the form: \(\sf\Large (\color{purple}{x_1},\color{red}{y_1})~and~(\color{green}{x_2},\color{blue}{y_2})\)

OpenStudy (nonoww):

-3/1

OpenStudy (nonoww):

right?

TheSmartOne (thesmartone):

Correct, and that can be simplified to just -3 now, point-slope form is \(y – y_1 = m(x – x_1)\) and m is the slope and you have to substitute one of the two points in. It doesn't matter which point

OpenStudy (nonoww):

im confused

TheSmartOne (thesmartone):

What is the slope?

OpenStudy (nonoww):

-3/1

TheSmartOne (thesmartone):

replace m with -3 \(y – y_1 = m(x – x_1)\)

OpenStudy (nonoww):

what about x1 and y 1

TheSmartOne (thesmartone):

you can use any point you want to use (1, 3) or (0, 6)

OpenStudy (nonoww):

what about x and y

TheSmartOne (thesmartone):

\( (x_1,y_1)\) \((~1, ~3~)\) \((~0,~6~)\) choose one point

OpenStudy (queenie):

leave x and y alone...sub in numbers for x1 and y1....and m

OpenStudy (precal):

you could also use y=mx+b b is your y intercept and you were given this point note: (0,6) b is 6 y=-3x+6 the slope is your m value and b is your y intercept